OpenStudy (anonymous):

Verify that the line with equation r=2i +4j +k +t(-4i+4j-5k) lies wholly in the plane with equation 3x-2y+4z=2

OpenStudy (perl):

wholly?

OpenStudy (anonymous):

Yeah, Wholly or Fully

OpenStudy (perl):

is there some theorem we can use?

OpenStudy (experimentx):

can you give me a point on vector r

OpenStudy (anonymous):

That's all it gives.

OpenStudy (dumbcow):

the vector normal to plane is <3,-2,4> if you can show that this is perpendicular to line vector then they are going same direction finally you have to show they then go through the same point look at t=0 r = 2i+4j+k, which is point (2,4,1) the plane also goes through this point --> 3*2 -2*4+4*1 = 2

OpenStudy (anonymous):

Why does t=0?

OpenStudy (experimentx):

isn't r = 2i+4j+k supposed to be position vector passing through 0,0,0 ?

OpenStudy (dumbcow):

to show 2 vectors are perpendicular, the dot product will be 0 <-4,4,-5> * <3,-2,4> = -40 ?? hmm is that maybe a positive 5?

OpenStudy (anonymous):

No. It's negative... I thought it would be too... so that it could be 2=2?

OpenStudy (dumbcow):

if the line is wholly in the plane, then wouldn't all the points for all t be on the plane?

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

@dumbcow isn't r supposed to be position vector passing through the origin??

OpenStudy (dumbcow):

yeah i guess..oh i am interpreting it like its a line equation

OpenStudy (dumbcow):

well would you agree, you could say x = 2-4t y=4+4t z=1-5t

OpenStudy (experimentx):

if it is then ... the given plane never passes through the origin and position vector passes through the origin moreover, when i studied 3d geometry, i studied line as (x-x1)/l = (y-y1)/n = ... k

OpenStudy (experimentx):